OpenStudy (anonymous):

Tommy Hayden decided to purchase a new moped on an installment loan. The moped was $1,500.00. He agreed to pay $150 a month for 12 months. What is the finance charge in dollars? The finance charge is $ i need help finding the finance charge.

OpenStudy (anonymous):

total cost=$150*12month=$1800 original cost of the mop is $1500 finance charge=1800-1500=300

OpenStudy (anonymous):

Thanks you! @eazykang